what is the sum of numbers between 0 and 270 which are divisible by 3
Answer by htmentor(1343)

what is the sum of numbers between 0 and 270 which are divisible by 3
=================================
The numbers form an arithmetic sequence: 0,3,6,9,...
The common difference is 3 and the 1st term is 0.
The sum of the 1st n terms of an arithmetic sequence is
S(n) = n/2(a_1 + a_n)
The number of terms, n = 91
So the sum is S(91) = 45.5*(0+270) = 12285